What is the standard form of equation of a line with x intercept of -3 and y intercept of 2/5?
Found 2 solutions by stanbon, checkley71:
Answer by stanbon(48558)

What is the standard form of equation of a line with x intercept of -3 and y intercept of 2/5?
--------------
The line passes through (0,2/5) and (-3,0)
slope = [2/5]/3=2/15
EQUATION: y=(2/15)x+(2/5)
To get the standard form multiply through by 15 to get:
15y=2x+6
2x-15y=-6
------------------
Cheers,
Stan H.

Answer by checkley71(8403)

THE 2 POINTS YOU HAVE ARE (-3,0) & (0,2/5) THUS THE SLOPE IS:
(2/5-0)/0+3
(2/5)/3
2/5*1/3
2/15 IS THE SLOPE (m) WITH A Y INTERCEPT OF 2/5. THUS THE EQUATION IS
Y=2/15X+2/5